OpenStudy (anonymous):

4 3/8 = k - 7 1/8 k = ?

OpenStudy (anonymous):

okay u have to get k by itself

OpenStudy (anonymous):

so in order to do that you have to do the inverse operation to both sides

OpenStudy (anonymous):

since its - 7 1/8 you have to add 7 1/8

OpenStudy (hyroko):

add 7&1/8 to both sides. Voila. you got k

OpenStudy (anonymous):

so cancel out the side with k and add 7 1/8 to 4 3/8

OpenStudy (hyroko):

4+3/8 +7+1/8 = 11 +1/2 (simplified from 4/8)

OpenStudy (anonymous):

since the denominators are the same add the numerators and whole numbers

OpenStudy (anonymous):

its 11 1/2

OpenStudy (anonymous):

thats k

OpenStudy (anonymous):

so ur whole numbers 4+7=11

OpenStudy (anonymous):

then numerators 1=3=4 put it over 8 and you have 4/8 simplified 4/8=1/2

OpenStudy (anonymous):

1/2 plus the whole numbers= 11 1/2

OpenStudy (anonymous):

1=3=4? are you on drugs?

OpenStudy (anonymous):

no it was simply a typo